# Envelopes and evolutes

### v1

@@ Abstract.
We define envelopes of families of linear spaces in projective space. In order to define evolutes we impose a notion of perpendicularity, which allows us to consider the normal spaces to a projective variety. The evolute of a projective hypersurface is the envelope of the family of normal spaces. For a variety of dimension $r$ in $n$-space, the evolute is defined as the (nr)(n-r)th “iterated” cuspidal locus of the map from the total space of the normal spaces to projective space. Thus the envelope can be interpreted as a (nr)(n-r)th order Thom–Boardman singularity. Further higher order Thom–Boardman singularities correspond, for a curve in the plane or in 3-space, to classical objects like the vertices of the curve; for a surface in 3-space, they give the cuspidal curve – and its cusps – on the evolute. Using known formulas for Thom polynomials we are able to verify and generalize many of Salmon’s formulas and find new ones.
@@ 1. Introduction
In this paper we consider projective complex varieties of any dimension and define the analogue of generalized evolutes of real curves. The evolutes are defined as “iterated cuspidal loci” of the envelope of the family of normal spaces to the given variety. These loci are Thom–Boardman singularities of the map from the total space of the family of normal spaces to the projective space, and hence their classes can be computed from known Thom polynomials. In Section 2 we define envelopes of families of linear spaces and determine the cycle class of their singular (cuspidal) loci. In Section 3 we introduce a notion of perpendicularity which allows us to define normal spaces to a projective variety. We can therefore consider the family of normal spaces to the given variety and define evolutes. In Section 4, we prove numerical formulas for the envelope of the normal spaces to curves in arbitrary dimensional projective space. We show that our formulas agrees with those of Salmon in the case of curves in 3-space. Section 5 computes various cycle classes associated to the evolute of a surface, and, in the case of a surface in 3-space, the degree of its evolute, the degree of the cuspidal curve of the evolute, and the number of cusps of the cuspidal curve. Section 6 treats osculating developables of curves in the context of envelopes.
@@
The author would also like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme “New equivariant methods in algebraic and differential geometry” in March 2024, where a part of the work on this paper was undertaken. This work was supported by EPSRC grant EP/R014604/1. Particular thanks go to László Fehér and Andrzej Weber for helpful discussions.
@@ 2. Envelopes of families of linear spaces
We fix a $\mathbb{C}$-vector space $V$ of dimension $n+1$. Let $X$ be a projective nonsingular variety of dimension $r$, and $f\colon X\to\mathbb{P}(V)$ a morphism. Assume $\mathcal{F}$ is a rank nr+1n-r+1 locally free sheaf on $X$ and $V_{X}\to\mathcal{F}$ a surjective map. Then $\mathcal{F}$ gives rise to a family of linear (nr)(n-r)-spaces
@@
where $\psi$ is the composition of the inclusion $\mathbb{P}(\mathcal{F})\subset X\times\mathbb{P}(V)$ with projection on the second factor $X\times\mathbb{P}(V)\to\mathbb{P}(V)$.
@@
The classes of these loci are given by Thom polynomials in the Chern classes $\overline{c}_{i}:=c_{i}(\psi^{*}T_{\mathbb{P}(V)}-T_{\mathbb{P}(\mathcal{F})})$. Recall that the Thom polynomial for $\Sigma^{1}$ is $\overline{c}_{1}$, that for $\Sigma^{1,1}$ is $\overline{c}_{1}^{2}+\overline{c}_{2}$ [Ronga], and that for $\Sigma^{1,1,1}$ is $\overline{c}_{1}^{3}+3\overline{c}_{1}\overline{c}_{2}+2\overline{c}_{3}$ [Rim]. (The Thom polynomial for $\Sigma^{1,1,1,1}$ was computed by Gaffney, Porteous, and Ronga [Gaf]*Thm. 2.2, p. 407:
@@
Since $\psi^{*}\mathcal{O}_{\mathbb{P}(V)}(1)=\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1)$, the first equality follows. For the other three, consider the two exact sequences
and
$0\to\Omega_{\mathbb{P}(\mathcal{F})/X}^{1}\otimes\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1)\to\mathcal{P}^{1}_{\mathbb{P}(\mathcal{F})/X}(1)=\pi^{*}\mathcal{F}\to\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1)\to 0.$
This gives
Together with the fact
$c_{k}(\pi^{*}\mathcal{F}\otimes\mathcal{O}_{\mathbb{P}(\mathcal{F})}(-1))=\sum_{i=0}^{k}(-1)^{k-i}\binom{n-r+1-i}{k-i}c_{i}(\pi^{*}\mathcal{F})c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))^{k-i},$
we conclude that $c_{i}=c_{i}(T_{\mathbb{P}(\mathcal{F})})=(-1)^{i}c_{i}(\Omega_{\mathbb{P}(\mathcal{F})}^{1})$ is as stated in the lemma. ∎
@@
The ramification locus of the restriction of the map of sheaves to $\Sigma^{1}$ is the second order Thom–Boardman singularity locus $\Sigma^{1,1}$. Denote by $C_{\mathcal{F}}$ the image by $\psi$ of this ramification locus, i.e., $C_{\mathcal{F}}$ is the branch locus of $\psi|_{E_{\mathcal{F}}}$. In case that $X$ is a curve in 3-space, $C_{\mathcal{F}}$ is the cuspidal edge of the developable surface $E_{\mathcal{F}}$.
@@
$C_{\mathcal{F}}]=\psi_{*}\bigl{(}\pi^{*}(2c_{1}(\Omega^{1}_{X})^{2}+3c_{1}(\Omega^{1}_{X})c_{1}(\mathcal{F})+2c_{1}(\mathcal{F})^{2}-c_{2}(\Omega^{1}_{X})-c_{2}(\mathcal{F}))+\\ (3r-1)\pi^{*}c_{1}(\mathcal{F})c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))+3r\pi^{*}c_{1}(\Omega^{1}_{X})c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))+\frac{3r^{2}-r}{2}c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))^{2}\bigr{)}\cap[\mathbb{P}(V)]$
@@
The class of the cuspidal locus $\kappa_{\mathcal{F}}$ of the cuspidal edge is
$[\kappa_{\mathcal{F}}]=\psi_{*}\bigl{(}(17\binom{r}{3}+12\binom{r}{2}+r)c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))^{3}+(17\binom{r}{2}+6r)c_{1}(\Omega^{1}_{X})c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))^{2}\\ \textstyle+(17\binom{r}{2}+r+2)c_{1}(\mathcal{F})c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))^{2}+11rc_{1}(\Omega^{1}_{X})^{2}c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))\\ +(17r-5)c_{1}(\Omega^{1}_{X})c_{1}(\mathcal{F})c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))+(11r-7)c_{1}(\mathcal{F})^{2}c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))\\ -5rc_{2}(\Omega^{1}_{X})c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))-(5r-4)c_{2}(\mathcal{F})c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))+6c_{1}(\Omega^{1}_{X})^{3}\\ +11c_{1}(\Omega^{1}_{X})^{2}c_{1}(\mathcal{F})+11c_{1}(\Omega^{1}_{X})c_{1}(\mathcal{F})^{2}+6c_{1}(\mathcal{F})^{3}-7c_{1}(\Omega^{1}_{X})c_{2}(\Omega^{1}_{X})\\ -7c_{2}(\Omega^{1}_{X})c_{1}(\mathcal{F})-5c_{1}(\Omega^{1}_{X})c_{2}(\mathcal{F})-7c_{1}(\mathcal{F})c_{2}(\mathcal{F})+2c_{3}(\Omega^{1}_{X})+2c_{3}(\mathcal{F})\bigr{)}\cap[\mathbb{P}(V)].$
@@
$\overline{c}_{1}=rc_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))+\pi^{*}(c_{1}(\Omega^{1}_{X})+c_{1}(\mathcal{F}))\\ \textstyle\overline{c}_{2}=\binom{r}{2}c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))^{2}+\pi^{*}(rc_{1}(\Omega^{1}_{X})+(r-1)c_{1}(\mathcal{F}))c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))\\ +\pi^{*}(c_{1}(\Omega^{1}_{X})^{2}+c_{1}(\Omega^{1}_{X})c_{1}(\mathcal{F})+c_{1}(\mathcal{F})^{2}-c_{2}(\Omega^{1}_{X})-c_{2}(\mathcal{F}))\\ \textstyle\overline{c}_{3}=\binom{r}{3}c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))^{3}+\pi^{*}(\binom{r}{2}c_{1}(\Omega_{X}^{1})+\binom{r-1}{2}c_{1}(\mathcal{F}))c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))^{2}\\ +\pi^{*}\bigl{(}rc_{1}(\Omega_{X}^{1})^{2}+(r-1)c_{1}(\Omega_{X}^{1})c_{1}(\mathcal{F})+(r-2)c_{1}(\mathcal{F})^{2}-rc_{2}(\Omega_{X}^{1})-(r-2)c_{2}(\mathcal{F})\bigr{)}c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))\\ +\pi^{*}\bigl{(}c_{3}(\mathcal{F})+c_{3}(\Omega_{X}^{1})+c_{1}(\Omega_{X}^{1})c_{2}(\mathcal{F})-2c_{1}(\Omega_{X}^{1})c_{2}(\Omega_{X}^{1})-2c_{1}(\Omega_{X}^{1})c_{2}(\mathcal{F})\\ -2c_{2}(\Omega_{X}^{1})c_{1}(\mathcal{F})-2c_{1}(\mathcal{F})c_{2}(\mathcal{F})+c_{1}(\Omega_{X}^{1})^{3}+c_{1}(\Omega_{X}^{1})^{2}c_{1}(\mathcal{F})+c_{1}(\Omega_{X}^{1})c_{1}(\mathcal{F})^{2}+c_{1}(\mathcal{F})^{3}\bigr{)},$
from which the proposition follows. ∎
@@ 3. Evolutes
In order to define the normal spaces to a projective varietiy, we need to equip $\mathbb{P}(V)$ with a “Euclidean structure”, by which we mean a notion of perpendicularity between linear spaces. Let $H_{\infty}:=\mathbb{P}(V^{\prime})$, where $V\to V^{\prime}$ is an $n$-quotient, be a hyperplane and $Q_{\infty}\subset H_{\infty}$ a non-singular quadric. If $L\subset H_{\infty}$ is an $i$-dimensional linear space, we let $L^{\perp}\subset H_{\infty}$ denote the (n2i)(n-2-i)-dimensional linear space which is polar to $L$ with respect to $Q_{\infty}$ [reci]*Sec. 4. If $L,L^{\prime}\subset\mathbb{P}(V)$ are linear spaces, we let $\langle L,L^{\prime}\rangle$ denote their linear span.
Assume that $X\subset\mathbb{P}(V)$ is a nonsingular variety of dimension $r$ and that $f(X)$ is in general position with respect to $H_{\infty}$ and $Q_{\infty}$ [reci]*Prop. 4.1, p. 245. Set $\mathcal{K}^{1}:=\operatorname{Ker}(V_{X}\to\mathcal{P}^{1}_{X}(1))$. Then $\mathcal{E}:=(\mathcal{K}^{1})^{\vee}\oplus\mathcal{O}_{X}(1)$ is the Euclidean normal bundle to $X$, and there is a natural surjective map $V_{X}\to\mathcal{E}$ (using the isomorphism $(V^{\prime})^{\vee}\cong V^{\prime}$ induced by the quadric $Q_{\infty}$).
The fibers of $\pi\colon\mathbb{P}(\mathcal{E})\to X$ are the (nr)(n-r)-spaces perpendicular to the tangent spaces (at the point of tangency) of $X$: if $P\in X$, then $\psi(\mathbb{P}(\mathcal{E})_{P})=\langle P,(T_{f(P)}\cap H_{\infty})^{\perp}\rangle$ (see [reci]*Sec. 4). Let $E_{\mathcal{E}}$ denote the envelope of the family $\mathbb{P}(\mathcal{E})\subset X\times\mathbb{P}(V)\to\mathbb{P}(V)$ of normal (nr)(n-r)-spaces to $X$, i.e., $E_{\mathcal{E}}$ is the branch locus of the composed map $\psi\colon\mathbb{P}(\mathcal{E})\to\mathbb{P}(V)$. We define the evolute of $X$ to be the image by $\psi$ of the (nr)(n-r)th order singularity locus $\Sigma^{1,\dots,1}$, where the number of 1’s is nrn-r. In particular, if $X$ is a hypersurface, then its evolute is equal to the envelope of its normal lines.
@@
This follows from Proposition 2.3, with $\mathcal{F}=\mathcal{E}=(\mathcal{K}^{1})^{\vee}\oplus\mathcal{O}_{X}(1)$, and using $c_{1}(\Omega_{X}^{1})=c_{1}(\mathcal{O}_{X}(d-n-1))$ and $\mathcal{K}^{1}=\mathcal{O}_{X}(-d+1)$.
Note that, for $d=2$, we get $\deg E_{\mathcal{E}}=6(n-1)$, which gives the very classical formulas for the degree of the evolute of a plane conic and of a quadric surface.
@@
If $f:X\to\mathbb{P}(V)$ is a morphism where $X$ is not necessarily nonsingular and $f$ is not necessarily an embedding, we may replace $X$ by its Nash transform $\nu:\widetilde{X}\to X$ and $V_{X}\to\mathcal{P}^{1}_{X}(1)$ by the Nash quotient $V_{\widetilde{X}}\to\mathcal{P}^{1}$ [reci]*Sec. 3.

### v2

@@ Abstract.
We define envelopes of families of linear spaces in projective space. In order to define evolutes we impose a notion of perpendicularity, which allows us to consider the normal spaces to a projective variety. The evolute of a projective hypersurface is the envelope of the family of normal lines. For a variety of dimension $r$ in $n$-space, the evolute is defined as the (nr)(n-r)th “iterated” cuspidal locus of the map from the total space of the normal spaces to projective space. Thus the envelope can be interpreted as a (nr)(n-r)th order Thom–Boardman singularity. Further higher order Thom–Boardman singularities correspond, for a curve in the plane or in 3-space, to classical objects like the vertices of the curve; for a surface in 3-space, they give the cuspidal curve – and its cusps – on the evolute. Using known formulas for Thom polynomials we are able to verify and generalize many of Salmon’s formulas and find new ones.
@@ 1. Introduction
In this paper we consider projective complex varieties of any dimension and define the analogue of generalized evolutes of real curves. The evolutes are defined as “iterated cuspidal loci” of the envelope of the family of normal spaces to the given variety. These loci are Thom–Boardman singularities of the map from the total space of the family of normal spaces to the projective space, and hence their classes can be computed from known Thom polynomials. In Section 2 we define envelopes of families of linear spaces and determine the cycle class of their singular (cuspidal) loci. In Section 3 we introduce a notion of perpendicularity which allows us to define normal spaces to a projective variety. We can therefore consider the family of normal spaces to the given variety and define evolutes. In Section 4, we prove numerical formulas for the envelope of the normal spaces to curves in arbitrary dimensional projective space. We show that our formulas agree with those of Salmon in the case of curves in 3-space. Section 5 computes various cycle classes associated to the evolute of a surface, and, in the case of a surface in 3-space, the degree of its evolute, the degree of the cuspidal curve of the evolute, and the number of cusps of the cuspidal curve. Section 6 treats osculating developables of curves in the context of envelopes.
@@
The author would also like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme “New equivariant methods in algebraic and differential geometry” in March 2024, where a part of the work on this paper was undertaken. This work was supported by EPSRC grant EP/R014604/1. Particular thanks go to László Fehér and Andrzej Weber for helpful discussions. Finally, the author thanks the referee for very useful comments.
@@ 2. Envelopes of families of linear spaces
We fix a $\mathbb{C}$-vector space $V$ of dimension $n+1$ and consider the corresponding projective $n$-dimensional space $\mathbb{P}(V)$. (We use the Grothendieck convention, so that points in $\mathbb{P}(V)$ are surjections $V\to\mathbb{C}$.) Let $X$ be a projective nonsingular variety of dimension $r$, and $f\colon X\to\mathbb{P}(V)$ a morphism. Assume $\mathcal{F}$ is a rank nr+1n-r+1 locally free sheaf on $X$ and $V_{X}\to\mathcal{F}$ a surjective map. Then $\mathcal{F}$ gives rise to a family of linear (nr)(n-r)-spaces
@@
where $\psi$ is the composition of the inclusion $\mathbb{P}(\mathcal{F})\subset X\times\mathbb{P}(V)$ with the projection on the second factor $X\times\mathbb{P}(V)\to\mathbb{P}(V)$. Note that $\dim\mathbb{P}(\mathcal{F})=\dim\mathbb{P}(V)$.
@@
The classes of these loci are given by Thom polynomials in the Chern classes $\overline{c}_{i}:=c_{i}(\psi^{*}T_{\mathbb{P}(V)}-T_{\mathbb{P}(\mathcal{F})})$, provided the map $\psi$ is suitably generic, in particular provided the loci $\Sigma^{1}$, $\Sigma^{1,1}$, and $\Sigma^{1,1,1}$ have the expected codimension. In what follows, we shall always assume this holds.
Recall that the Thom polynomial for $\Sigma^{1}$ is $\overline{c}_{1}$, that for $\Sigma^{1,1}$ is $\overline{c}_{1}^{2}+\overline{c}_{2}$ [Ronga], and that for $\Sigma^{1,1,1}$ is $\overline{c}_{1}^{3}+3\overline{c}_{1}\overline{c}_{2}+2\overline{c}_{3}$ [Rim]. (The Thom polynomial for $\Sigma^{1,1,1,1}$ was computed by Gaffney, Porteous, and Ronga [Gaf]*Thm. 2.2, p. 407:
@@
Lemma 2.2.
We have
Proof.
The exact sequence
gives $c(\Omega_{\mathbb{P}(\mathcal{F})}^{1})=\pi^{*}c(\Omega^{1}_{X})c(\Omega_{\mathbb{P}(\mathcal{F})/X}^{1})$, and the sequence
$0\to\Omega_{\mathbb{P}(\mathcal{F})/X}^{1}\otimes\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1)\to\mathcal{P}^{1}_{\mathbb{P}(\mathcal{F})/X}(1)=\pi^{*}\mathcal{F}\to\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1)\to 0$
gives $c(\Omega_{\mathbb{P}(\mathcal{F})/X}^{1})=c(\pi^{*}\mathcal{F}\otimes\mathcal{O}_{\mathbb{P}(\mathcal{F})}(-1))$. ∎
@@
Since $\psi^{*}\mathcal{O}_{\mathbb{P}(V)}(1)=\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1)$, the first equality follows.
For the other three, we use Lemma 2.2 and the fact that
$c_{k}(\pi^{*}\mathcal{F}\otimes\mathcal{O}_{\mathbb{P}(\mathcal{F})}(-1))=\sum_{i=0}^{k}(-1)^{k-i}\binom{n-r+1-i}{k-i}c_{i}(\pi^{*}\mathcal{F})c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))^{k-i}.$
It follows that $c_{i}=c_{i}(T_{\mathbb{P}(\mathcal{F})})=(-1)^{i}c_{i}(\Omega_{\mathbb{P}(\mathcal{F})}^{1})$ is as stated in the lemma. ∎
@@
The ramification locus of the restriction of the map of sheaves to $\Sigma^{1}$ is the second order Thom–Boardman singularity locus $\Sigma^{1,1}$. Denote by $C_{\mathcal{F}}$ the image by $\psi$ of this ramification locus, i.e., $C_{\mathcal{F}}:=\psi_{*}(\Sigma^{1,1})$ is the branch locus of $\psi|_{E_{\mathcal{F}}}$. In case that $X$ is a curve in 3-space, $C_{\mathcal{F}}$ is the cuspidal edge of the developable surface $E_{\mathcal{F}}$.
@@
$C_{\mathcal{F}}]=\psi_{*}\bigl(\pi^{*}(2c_{1}(\Omega^{1}_{X})^{2}+3c_{1}(\Omega^{1}_{X})c_{1}(\mathcal{F})+2c_{1}(\mathcal{F})^{2}-c_{2}(\Omega^{1}_{X})-c_{2}(\mathcal{F}))\\ +(3r-1)\pi^{*}c_{1}(\mathcal{F})c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))+3r\pi^{*}c_{1}(\Omega^{1}_{X})c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))\\ +\frac{3r^{2}-r}{2}c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))^{2}\bigr)\cap[\mathbb{P}(V)],$
@@
Set $h:=c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))=\psi^{*}c_{1}(\mathcal{O}_{\mathbb{P}(V)}(1)$, $w_{i}:=\pi^{*}c_{i}(\Omega^{1}_{X})$, and $f_{i}:=\pi^{*}c_{i}(\mathcal{F})$. The class of the cuspidal locus $\kappa_{\mathcal{F}}:=\psi_{*}(\Sigma^{1,1,1})$ of the cuspidal edge $C_{\mathcal{F}}$ is
$[\kappa_{\mathcal{F}}]=\psi_{*}\bigl((17\binom{r}{3}+12\binom{r}{2}+r)h^{3}+\bigl((17\binom{r}{2}+6r)w_{1}\textstyle+(17\binom{r}{2}+r+2)f_{1}\bigr)h^{2}\\ +\bigl(11rw_{1}^{2}+(17r-5)w_{1}f_{1}+(11r-7)f_{1}^{2}-5rw_{2}-(5r-4)f_{2}\bigr)h\\ +6w_{1}^{3}+11w_{1}^{2}f_{1}+11w_{1}f_{1}^{2}+6f_{1}^{3}-7w_{1}w_{2}-7w_{2}f_{1}-5w_{1}f_{2}-7f_{1}f_{2}+2w_{3}+2f_{3}\bigr)\cap[\mathbb{P}(V)].$
@@
$\overline{c}_{1}=rc_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))+\pi^{*}(c_{1}(\Omega^{1}_{X})+c_{1}(\mathcal{F}))\\ \textstyle\overline{c}_{2}=\binom{r}{2}c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))^{2}+\pi^{*}(rc_{1}(\Omega^{1}_{X})+(r-1)c_{1}(\mathcal{F}))c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))\\ +\pi^{*}(c_{1}(\Omega^{1}_{X})^{2}+c_{1}(\Omega^{1}_{X})c_{1}(\mathcal{F})+c_{1}(\mathcal{F})^{2}-c_{2}(\Omega^{1}_{X})-c_{2}(\mathcal{F}))\\ \textstyle\overline{c}_{3}=\binom{r}{3}c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))^{3}+\pi^{*}(\binom{r}{2}c_{1}(\Omega_{X}^{1})+\binom{r-1}{2}c_{1}(\mathcal{F}))c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))^{2}\\ +\pi^{*}\bigl(rc_{1}(\Omega_{X}^{1})^{2}+(r-1)c_{1}(\Omega_{X}^{1})c_{1}(\mathcal{F})+(r-2)c_{1}(\mathcal{F})^{2}-rc_{2}(\Omega_{X}^{1})\\ -(r-2)c_{2}(\mathcal{F})\bigr)c_{1}(\mathcal{O}_{\mathbb{P}(\mathcal{F})}(1))+\pi^{*}\bigl(c_{3}(\mathcal{F})+c_{3}(\Omega_{X}^{1})+c_{1}(\Omega_{X}^{1})c_{2}(\mathcal{F})\\ -2c_{1}(\Omega_{X}^{1})c_{2}(\Omega_{X}^{1})-2c_{1}(\Omega_{X}^{1})c_{2}(\mathcal{F})\-2c_{2}(\Omega_{X}^{1})c_{1}(\mathcal{F})-2c_{1}(\mathcal{F})c_{2}(\mathcal{F})\\ +c_{1}(\Omega_{X}^{1})^{3}+c_{1}(\Omega_{X}^{1})^{2}c_{1}(\mathcal{F})+c_{1}(\Omega_{X}^{1})c_{1}(\mathcal{F})^{2}+c_{1}(\mathcal{F})^{3}\bigr),$
from which the proposition follows. Note that the expression is polynomial in $r$ and independent of $n$. ∎
@@ 3. Evolutes
In order to define the normal spaces to a projective variety, we need to equip $\mathbb{P}(V)$ with a “Euclidean structure”, by which we mean a notion of perpendicularity between linear spaces. Let $V^{\prime}$ be a vector space of dimension $n$ and $V\to V^{\prime}$ a surjection, so that $\mathbb{P}(V^{\prime})\subset\mathbb{P}(V)$ is a hyperplane. Set $H_{\infty}:=\mathbb{P}(V^{\prime})$, the hyperplane at infinity. Let $Q_{\infty}\subset H_{\infty}$ be a non-singular quadric hypersurface. If $L\subset H_{\infty}$ is an $i$-dimensional linear space, we let $L^{\perp}\subset H_{\infty}$ denote the (n2i)(n-2-i)-dimensional linear space which is polar to $L$ with respect to $Q_{\infty}$ [reci]*Sec. 4, p. 244. If $L,L^{\prime}\subset\mathbb{P}(V)$ are linear spaces, we let $\langle L,L^{\prime}\rangle$ denote their linear span.
Assume that $X\subset\mathbb{P}(V)$ is a nonsingular variety of dimension $r$ and that $X$ is in general position with respect to $H_{\infty}$ and $Q_{\infty}$, i.e., $H_{\infty}$ and $X$, and $Q_{\infty}$ and $X\cap H_{\infty}$, intersect transversally. Let $\mathcal{P}^{1}_{X}(1)$ denote the sheaf of principal parts of $\mathcal{O}_{X}(1)$ of order 1, which is a vector bundle of rank $r+1$. The first jet map $V_{X}\to\mathcal{P}^{1}_{X}(1)$ is surjective, and the fibers $T_{P}:=\mathbb{P}(\mathcal{P}^{1}_{X}(1))_{P}\subset\mathbb{P}(V)$ are the projective tangent spaces to $X$. Set $\mathcal{K}^{1}:=\operatorname{Ker}(V_{X}\to\mathcal{P}^{1}_{X}(1))$, so that $\mathcal{K}^{1}$ is the conormal bundle of $X$ in $\mathbb{P}(V)$ twisted by $\mathcal{O}_{X}(1)$. Then $\mathcal{E}:=(\mathcal{K}^{1})^{\vee}\oplus\mathcal{O}_{X}(1)$ is the Euclidean normal bundle to $X$, and there is a natural surjective map $V_{X}\to\mathcal{E}$ (using the isomorphism $(V^{\prime})^{\vee}\cong V^{\prime}$ induced by the quadric $Q_{\infty}$) [reci]*Prop. 4.1, p. 245.
The fibers $N_{P}:=\mathbb{P}(\mathcal{E})_{P}\subset\mathbb{P}(V)$ are the normal spaces to $X$, i.e., the (nr)(n-r)-spaces perpendicular to the tangent spaces $T_{P}$. Indeed, $N_{P}=\langle P,(T_{P}\cap H_{\infty})^{\perp}\rangle$ (see [reci]*Sec. 4).
Let $E_{\mathcal{E}}$ denote the envelope of the family $\mathbb{P}(\mathcal{E})\subset X\times\mathbb{P}(V)\to\mathbb{P}(V)$ of normal (nr)(n-r)-spaces to $X$, i.e., $E_{\mathcal{E}}$ is the branch locus of the composed map $\psi\colon\mathbb{P}(\mathcal{E})\to\mathbb{P}(V)$. We define the evolute of $X$ to be the image by $\psi$ of the (nr)(n-r)th order singularity locus $\Sigma^{1,\dots,1}$, where the number of 1’s is nrn-r. In particular, if $X$ is a hypersurface, then its evolute is equal to the envelope of its normal lines.
@@
This follows from Proposition 2.4, with $\mathcal{F}=\mathcal{E}=(\mathcal{K}^{1})^{\vee}\oplus\mathcal{O}_{X}(1)$. Since $X$ is a hypersurface of degree $d$, we have $c_{1}(\Omega_{X}^{1})=c_{1}(\mathcal{O}_{X}(d-n-1))$ and $\mathcal{K}^{1}=\mathcal{O}_{X}(-d)\otimes\mathcal{O}_{X}(1)=\mathcal{O}_{X}(-d+1)$.
Note that, for $d=2$, we get $\deg E_{\mathcal{E}}=6(n-1)$, which gives the very classical formulas for the degree of the evolute of a plane conic ($n=2$) and of a quadric surface ($n=3$).
@@
If $f:X\to\mathbb{P}(V)$ is a morphism where $X$ is not necessarily nonsingular and $f$ is not necessarily an embedding, we may replace $X$ by its Nash transform $\nu:\widetilde{X}\to X$ and $V_{X}\to\mathcal{P}^{1}_{X}(1)$ by its Nash quotient $V_{\widetilde{X}}\to\mathcal{P}^{1}$ [reci]*Sec. 3. In this case, the genericity assumption is that $H_{\infty}$ intersects each Whitney stratum of $f(X)$ transversally, and that $Q_{\infty}$ intersects each Whitney stratum of $f(X)\cap H_{\infty}$ transversally [reci]*Sec. 4, p. 245.